PartnerObserving Log for 2007 May 10/11

Technical Observers: J. Hill, A. DiPaola, R. Speziali
Telescope Operator: E. Lundin
Instrument: LBC-Blue

Some more SDT-related observing during engineering time. Note that I'm paying little attention to the SDT queue. I'm only using convenient SDT targets to test various aspects of the telescope and instrument.

New clad fibers have been installed in the LBC-Blue hub to solve the leaking light problem in a permanent way.

New Dataprobe iP-2815's have been installed to improve LBC-Blue power management.

Lost first 1.6 hours to a problem with HBS (hydrostatic bearing system). The day guys left a wire disconnected while they were doing upgrades to pressure sensors.

Opened at 04:00 UT. Sky is crystal clear. T=10degC D=-8degC Calm.

Lost next 0.6 hours to the Left Mirror Support Compressor overheating. Switched to running both mirrors on the Right Compressor. In parallel to that, the Right Mirror Ventilation Counterflow Fan tripped its breaker.

5 biases

04:44 - 900 sec dark with dome open

05:04 - Focus near M81

05:16 - M81V - Seeing not very good at 2 arcsec. Note large temperature change (5degC) since last night and late opening.

05:25 - Focus at Lockman Block 3.

05:29 - Egami_Lockman_XMM_U_1 - Images improve to 1.5 arcsec in Uspec.

06:06 - Focus at Lockman Block 3 Dither 5.

06:09 - Egami_Lockman_XXM_U_2 - Images still about 1.5 arcsec. I think I see field-dependent astigmatism now that Emiliano has left the mountain. ??

06:50 - PG1323_UB

06:59:30 - LBC-Blue Filter Wheel #2: "Timeout error on stop". Turn camera off/on. 07:07:56 - still getting filter wheel errors after the restart "Motion Stop Return Error" on both wheels. Turn camera on/off again. Still errors Went to visit instrument in person. Both filter wheels are free to turn by hand. New fibers are NOT wrapped up in the filter wheel!

07:39 - Stopped the LBC program and ran "testfilters" on CMU. Both filter wheels work fine with testfilters.

07:48 - Restarted the LBC program and turned on the camera again.

07:58 - Vfastextra to focus at position of PG1323. Filter wheels work on the first move. Lost 1 hour to this unknown filter wheel problem.

08:01 - PG1323_Vgr

08:09 - IC_4182_U_photo

08:13 - Vfastextra at IC_4182

08:15 - IC_4182_V_photo - Seeing and/or focus is much improved. 0.7 arcsec probably dominated by tracking/guiding jitter.

08:20 - IC_4182_U_photo - again with better images

08:25 - IC_4182_U 10step - 0.8 arcsec U images at 1.4 airmasses -- much better! Guiding failed on dither 8.

09:03 - Vfastextra at IC_4182

09:05 - IC_4182_V 10step - Image jumped badly (arcsec) at 75 seconds into dither 7.

Stopping science for a pointing check.

09:47 - 1sec B exposure of pointing star (1075, 3194) Az=425 El=73

09:53 - 1sec B exposure of pointing star (1244, 3128) Az=392 El=79 (two stars?)

09:57 - 1sec B exposure of pointing star (1270, 3095) Az=384 El=73

10:08 - 1 sec B exposure of pointing star (866, 3186) Az=110 El=61

10:10 - 1 sec B exposure of pointing star (931, 3179) Az=102 El=62

All this pointing data is wrong since I still had an SX repointing in from last nights experiments. Aaarghhh.

Attempting to measure image quality as a function of primary mirror tilt (traded against decenter to correct coma).

10:16 - Bfastextra at the field of IC4954 El=71 20:04:48 +29:15

10:28 - B5sec Xg=2.70 Yg=0 Zg=0 RXg=0 RYg=-60

10:32 - Bfastextra .....

10:36 - B5sec with all the global offsets zeroed.

10:46 - Bfastextra .....

10:50 - B5sec with Xg=+4.01 RYg=-90

EL=72.6 Xt=2.375 Yt=-0.535 Zt=-0.530 RXt=-1.0 RYt=-60

10:58 - 1 sec B exposure of pointing star (1267, 3341) Az=31 El=55 Still with this large mirror repointing.......on purpose.

Checking the real pointing again.....

11:02 - 1 sec B exposure of pointing star (1027, 2982) Az=31 El=55 with all the global offsets zeroed.

11:07 - 1 sec B exposure of pointing star (986, 2962) Az=79 El=21 with all the global offsets zeroed.

11:13 - 1 sec B exposure of pointing star (989, 2975) Az=96 El=18 with all the global offsets zeroed.

11:19 - PG1633_UB

11:27 - PG1633_Vgr

11:34 - SkyFlatTest_V ==> 700 cts per sec at the position of PG1633 Az=250 El=43

11:36 - SkyFlatV_5 x1 (aborted)

11:38 - SkyFlatV_5 x10 (????)

11:40 - SkyFlatV_5 x10 ==> 20000, 24000, 27000, 30000, 36000

11:45 - SkyFlatVRot180_5 x1 Andrea caught me by fixing the feature that keeps the exposure time scaling. Now it resets to 1.0 for each new OB. Paused and reset it to x5.

11:49 - SkyFlatVRot180_5 x5 47000, 56000, saturated.

11:52 - SkyFlatU_5 x5

11:58 - SkyFlatURot180_5 x3 (The 4th one in the set saturated.)

12:05 - Close Dome T=9.5degC, Wind Calm

Moth season has officially arrived! Batten down the hatches.

12:15 0 minute shutter open dark at zenith. Dome lights are off, and twilight rising outside. Too much light leaking in from outside, so we didn't learn anything about leaking fibers. That experient will have to wait for a cloudy night.

25 biases

-- JohnHill - 11 May 2007
Topic revision: r7 - 21 Jul 2009, NormCushing
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback